### v1

# Hierarchical Relation-augmented Representation Generalization for Few-shot Action Recognition

Abstract.
Few-shot action recognition (FSAR) aims to recognize novel action categories with few exemplars. Existing methods typically learn frame-level representations independently for each video by designing various inter-frame temporal modeling strategies. However, they neglect explicit relation modeling between videos and tasks, thus failing to capture shared temporal patterns across videos and reuse temporal knowledge from historical tasks. In light of this, we propose HR2G-shot, a Hierarchical Relation-augmented Representation Generalization framework for FSAR, which unifies three types of relation modeling (inter-frame, inter-video, and inter-task) to learn task-specific temporal patterns from a holistic view. In addition to conducting inter-frame temporal interactions, we further devise two components to respectively explore inter-video and inter-task relationships: i) Inter-video Semantic Correlation (ISC) performs cross-video frame-level interactions in a fine-grained manner, thereby capturing task-specific query features and learning intra- and inter-class temporal correlations among support features; ii) Inter-task Knowledge Transfer (IKT) retrieves and aggregates relevant temporal knowledge from the bank, which stores diverse temporal patterns from historical tasks. Extensive experiments on five benchmarks show that HR2G-shot outperforms current top-leading FSAR methods. Full code will be released.
copyright: none
1. Introduction
During the last few years, action recognition (Feichtenhofer et al., 2019; Wang et al., 2021a; Xing et al., 2023a; Huang et al., 2023, 2024a; Cao et al., 2024; Li et al., 2024) has witnessed remarkable progress with the advances in deep learning. However, their strong performance heavily relies on a large amount of labeled training examples, which restricts the model’s scalability and generalizability in data scarcity scenarios. In contrast, human beings can easily learn new visual concepts with only a few supervisions. To enable the machine to acquire such ability, many studies have shifted their attention towards few-shot action recognition (FSAR) (Zhu and Yang, 2018; Cao et al., 2020; Wang et al., 2024c), which aims to learn novel (unseen) action classes using only a few annotated video samples after training on a set of base (seen) classes with abundant samples.
Current top-leading FSAR solutions typically adopt the metric-based meta-learning paradigm, wherein the model first maps the query videos and support videos into discriminative feature space, and then performs query-support video matching based on pre-defined or learned distance metrics. To achieve this, a prevalent of subsequent efforts delve into network designs for temporal representation learning by, e.g., temporal attention operations (Wang et al., 2022), elaborated spatio-temporal (Thatipelli et al., 2022) or temporal-channel interaction (Xing et al., 2023c), multi-layer feature fusion (Xing et al., 2023b), and auxiliary modal information (e.g., depth (Fu et al., 2020) and motion (Wang et al., 2023a)). Apart from investigating temporal representation learning, to perform proper temporal feature alignment when comparing query and support videos, recent approaches devise various temporal matching strategies, e.g., frame-level feature matching (Cao et al., 2020; Wang et al., 2024c), tuple-level feature matching (Perrett et al., 2021), frame-to-segment feature matching (Wu et al., 2022), or multi-level feature matching (Zheng et al., 2022).
Though achieving remarkable performance, these methods suffer from two limitations: First, they independently learn frame-level representations for each video, neglecting the relation modeling and holistic semantic patterns among video samples in the current task. Exploring inter-video relationships can not only learn subtle semantic discrepancies among different videos, but also capture shared similar temporal patterns (e.g., action styles and speed) from diverse video samples, particularly under data-limited conditions. Although some preliminary attempts (Xing et al., 2023c; Wang et al., 2022) have been made towards this, they capture sample-level correlations at the video level (global feature for each video), ignoring fine-grained temporal details in videos. Thus, these methods fail to capture shared temporal patterns across videos in each task.
Figure 1. Our main idea. Previous FSAR works only rely on inter-frame relation modeling to learn video representations, ignoring the relations between videos and tasks. In contrast, we unify three types of relation modeling (i.e., inter-frame, inter-video, inter-task) under one single framework, so as to capture task-specific temporal cues.
Second, existing methods directly learn meaningful video representation from one single task, lacking explicit modeling of task-relatedness (i.e., some tasks share similar action patterns or structures). Thus, they fail to reuse experience from historical tasks and excavate task-shared temporal knowledge, causing sub-optimal transferability and generalization on new tasks with only a few samples. In light of the above, our goal is to unify three types of relation modeling (i.e., inter-frame, inter-video, inter-task) under one single framework, so as to yield task-specific temporal features for each video via transferring knowledge at both video-level and task-level (as shown in Fig. 1).
In this vein, we develop a novel hierarchical framework for FSAR, namely HR2G-shot, which learns task-specific temporal cues from multiple perspectives (i.e., inter-frame, inter-video, inter-task) in a progressive manner. In addition to performing inter-frame temporal interactions, we further explicitly learn inter-video and inter-task relationships, so as to strengthen temporal pattern understanding for each video from a holistic view. Thus, we propose two feasible modules: i) Inter-video Semantic Correlation (ISC) aims to precisely capture inter-video temporal relationships within each task. Specifically, we perform cross-video frame-level interactions in a fine-grained manner to gain task-specific temporal features and reduce considerable computational costs compared with inter-video dense interaction (see Table 3). Moreover, to respect the difference between support and query videos, we customize the masked interaction strategy for support-support and query-support relations (i.e., support and query videos only aggregate semantic information from support videos). By this means, our model could interactively transfer temporal knowledge between videos to yield task-specific query features and learn intra- and inter-class correlation among support features. ii) Inter-task Knowledge Transfer (IKT) is devised to learn transferable temporal knowledge between tasks by exploring inter-task relationships. Concretely, we develop a temporal knowledge bank to store diverse temporal patterns from historical tasks. Then we make use of temporal prototypes, which summarize frame-level representations of each video, to retrieve and aggregate temporal knowledge related to new tasks from the bank. Integrated with such knowledge, our method yields more meaningful video representations, thus quickly adapting to new tasks with only a few samples. We conduct extensive experiments on five gold-standard datasets, and the results demonstrate that our HR2G-shot outperforms existing top-leading FSAR methods by a large margin.
Overall, the main contributions of this work are threefold:
  • We provide a hierarchical relation-augmented framework for FSAR, which unifies three types of relation modeling to learn task-specific temporal patterns.
  • To explore inter-video temporal relationships within each task, we propose Inter-video Semantic Correlation that conducts cross-video frame-level interactions in a fine-grained manner and customizes masked interaction strategy for support-support and query-support relations.
  • To learn transferable temporal knowledge between tasks, we design Inter-task Knowledge Transfer that retrieves and aggregates relevant temporal knowledge from the temporal knowledge bank, which stores diverse temporal patterns from historical tasks.
2. Related Work
Few-shot Image Classification. The goal of few-shot image classification (Fei-Fei et al., 2006) is to recognize unseen categories with limited annotated samples. Existing few-shot solutions can be broadly categorized into three groups: i) Augmentation-based methods typically generate auxiliary data samples by feature augmentation (Chen et al., 2019a, b; Albalak et al., 2024) or generative models (Li et al., 2020; Zhang et al., 2018), to increase the diversity of the support set; ii) Optimization-based methods (Jamal and Qi, 2019; Sun and Gao, 2024; Rusu et al., 2019; Rajeswaran et al., 2019; Finn et al., 2017) learn good initial model parameters, such that updating the initial parameters via a few gradient steps could adapt well to new tasks; iii) Metric-based methods (Yoon et al., 2019; Yang et al., 2024; Zhu and Koniusz, 2023; Cheng et al., 2023) (e.g., ProtoNet (Snell et al., 2017) and RelationNet (Sung et al., 2018)) first learn a common feature space for all classes, and then employ pre-defined or learned distance metrics to compare query and support samples.
Our work shares a similar spirit of metric-based methods (Snell et al., 2017; Vinyals et al., 2016), whereas we address more complex and challenging few-shot action recognition that requires long-range relation modeling between spatio-temporal locations. With respect to this, we further jointly learn inter-video and inter-task relationships to learn distinct temporal features. By unifying such three types of relation modeling under our framework, we can learn task-specific temporal cues for each video from a holistic view.
3.3)3.4)$\bm{X}_{\mathrm{s}},\bm{X}_{\mathrm{q}}$$\bm{F}_{\mathrm{s}},\bm{F}_{\mathrm{q}}$$\tilde{\bm{F}}_{\mathrm{s}},\tilde{\bm{F}}_{\mathrm{q}}$(Eq. 7)
Figure 2. The overview of HR2G-shot. (a) HR2G-shot unifies three types of relation modeling (i.e., inter-frame, inter-video, and inter-task) to learn discriminative temporal features. (b) Inter-video Semantic Correlation (ISC) conducts fine-grained cross-video interactions to learn inter-video relationships. (c) To explore inter-task relationships, we retrives and aggregates temporal knowledge from the bank, which maintains diverse temporal patterns from historical tasks.
Few-shot Action Recognition (FSAR). FSAR is gaining much attention due to its practical value in reducing manual annotations at a large scale required by traditional action recognition. Compared with few-shot image classification, FSAR is more challenging due to the rich spatial-temporal information in videos. Towards FSAR, most existing solutions (Wang et al., 2024a; Wu et al., 2024; Liu et al., 2022; Tang et al., 2023) belong to the meta-learning paradigm (Snell et al., 2017; Vinyals et al., 2016), wherein the model learns discriminative temporal features from base action categories, and generalizes learned temporal representations on novel action categories. Based on the meta-learning policy, a main group of efforts focus on temporal representation learning. Early works (Zhu and Yang, 2018, 2020) directly aggregate frame-level features to learn a global representation for each video. Though straightforward, these methods neglect complex temporal cues within one video, leading to suboptimal performance. To address this limitation, recent methods attempt to model temporal relations by, e.g., learning long-range temporal dependencies (Thatipelli et al., 2022; Wang et al., 2022), detailed cross-frame patch-level interactions (Zhang et al., 2023), exploiting low-level spatial features via feature architecture search (Xing et al., 2023b), or multi-modal feature fusion (Wanyan et al., 2023; Fu et al., 2020; Wang et al., 2021b; Huang et al., 2024c). Apart from investigating video feature representation, another line of work turns to designing various feature matching strategies, e.g., frame-level feature alignment (Wang et al., 2022), tuple-level feature alignment (Perrett et al., 2021), and even multi-level feature alignment (Zheng et al., 2022; Wu et al., 2024; Huang et al., 2024b). For example, OTAM (Cao et al., 2020) performs frame-level feature alignment between videos with the DTW algorithm, which takes the temporal ordering characteristics of videos into account. HyRSM (Wang et al., 2022) designs a novel bidirectional Mean Hausdorff metric to obtain temporal matching scores between support and query videos from the set matching perspective.
Despite promising results have been achieved, they typically learn temporal representations independently for each video, ignoring the relation modeling between video samples and tasks. As a result, they fail to gain a holistic semantic pattern understanding from multi-level perspectives. Though a few works (e.g., GgHM (Xing et al., 2023c) and HyRSM (Wang et al., 2022)) attempt to learn task-specific embeddings by considering inter-task relationships, they only focus on the correlations between samples at the video level, overlooking fine-grained temporal interactions among videos. In contrast, our HR2G-shot i) learns inter-video relationships at fine-grained frame level, and ii) gather temporal knowledge from historical tasks to achieve inter-task knowledge transfer, thereby capturing task-specific temporal features for each video and ensuring a sufficient understanding of holistic semantic patterns for each task.
3. Method
3.1. Problem Formulation
The objective of few-shot action recognition (FSAR) is to classify novel action categories with limited samples per class. Under the few-shot setting, the model generalizes learned knowledge on training set $\mathcal{D}_{base}$ to novel testing set $\mathcal{D}_{novel}$, where base classes $\mathcal{C}_{base}$ and novel classes $\mathcal{C}_{novel}$ are disjoint. Consistent with previous works (Cao et al., 2020; Wang et al., 2022; Tang et al., 2024), we formulate FSAR problem with episodic training and testing . Specifically, each $N$-way $K$-shot episode task involves sampled $K$ labeled videos from each of $N$ different action classes as the support set, and a portion of the remaining samples from $N$ classes as the query set. During inference, we randomly sample multiple episodic tasks from $\mathcal{D}_{novel}$ and present the average results over these tasks, in order to comprehensively evaluate the performance of the few-shot model.
3.2. Algorithm Overview
We introduce HR2G-shot, which hierarchically models multi-level relations (i.e., inter-frame, inter-video, and inter-task) to capture task-specific temporal patterns from a holistic view for FSAR. The overall architecture of HR2G-shot is shown in Fig. 2. Given an $N$-way $K$-shot episode task with $NK$ support videos and $L$ query videos, we employ CLIP visual encoder (Radford et al., 2021) to get frame-level features for support and query videos, i.e., $\bm{X}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$ and $\bm{X}_{\mathrm{q}}\in{\mathbb{R}}^{L\times T\times C}$, where $T$ is the length of sampled frames in each video. Note that we adopt Parameter-Efficient Fine-Tuning (PEFT) strategy to fine-tune the visual encoder with minimal trainable parameters as in (Xing et al., 2024; Qu et al., 2025; Yang et al., 2023). Next, following previous works (Wang et al., 2024b), we concatenate frame-level features and corresponding textual features along the temporal dimension and then feed into temporal Transformer (Wang et al., 2024b, 2023a) (i.e., Inter-frame Temporal Modeling). In this way, we perform inter-frame interactions and obtain enhanced support and query features. To compensate frame-level relation modeling from other semantic perspectives, we further design two feasible modules: i) Inter-video Semantic Correlation (ISC) (§3.3) learns inter-video temporal relationships within each task by fine-grained cross-video interactions; ii) Inter-task Knowledge Transfer (IKT) (§3.4) aggregates useful temporal knowledge from historical tasks to learn inter-task relationships. Finally, the obtained task-specific support and query features, i.e., $\bar{\bm{F}}_{\mathrm{s}}$ and $\bar{\bm{F}}_{\mathrm{q}}$, are fed into temporal matching metrics to get the class predictions.
3.3. Inter-video Semantic Correlation
Different from exploiting inter-frame interactions in each video, capturing inter-video relationships can strengthen task-specific semantic patterns for each video. To achieve this goal, we design Inter-video Semantic Correlation (ISC) module to conduct Fine-grained Inter-video Interaction and further learn intra- and inter-class correlation by Adaptive Interaction Mask (Fig. 3(a)) for support-query and support-support videos (Fig. 2(b)). Via Adaptive Interaction Mask, support and query videos aggregate features from different types of videos (i.e., support or query) based on their respective characteristics. Compared with previous works (Wang et al., 2022; Xing et al., 2023c), which only consider utilize global video features to perform interactions among support and query videos, we attempt to make use of temporal cues in the video to strengthen sufficient temporal pattern understanding among videos.
Fine-grained Inter-video Interaction. Based on the above ananlysis, how to perform fine-grained inter-video interactions is crucial for FSAR. A straightforward approach is to perform dense attention along temporal and sample dimensions simultaneously. Thus, given a task with $NK$ support videos and $L$ query videos, the computational complexity of such dense sample attention is $O((NK+L)^{2}T^{2})$, causing high computational costs. Besides, due to the difference between the temporal dimension and the sample dimension, it is not reasonable to treat them equivalently.
Thus, we decouple dense sample attention into inter-frame interaction and our designed Inter-video Semantic Correlation (ISC). The inter-frame interaction is achieved by previous Inter-frame Temporal Modeling, obtaining support and query temporal features, i.e., $\bm{F}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$ and $\bm{F}_{\mathrm{q}}\in{\mathbb{R}}^{L\times T\times C}$. Next, ISC concatenates support and query features, and reshapes these features in one task to ${\mathbb{R}}^{T\times(NK+L)\times C}$. Afterwards, multi-head self-attention along with sample attention matrix $\bm{A}\in{\mathbb{R}}^{(NK+L)\times(NK+L)}$ is adopted to conduct cross-video interactions by attending all frame features at the same temporal location, resulting in task-specific temporal futures for support and query videos:
(1) $\tilde{\bm{F}}=[\tilde{\bm{F}}_{\mathrm{s}},\tilde{\bm{F}}_{\mathrm{q}}]=\texttt{SelfAttn}(\texttt{Re}(\texttt{Cat}(\bm{F}_{\mathrm{s}},\bm{F}_{\mathrm{q}}))),\vspace{-2pt}$
where Re is the reshape operation, and Cat denotes the concatenation operation. SelfAttn represents the standard multi-head self-attention. As such, ISC not only captures task-specific temporal features via interactively transfer temporal knowledge between videos, but also reduces the computational complexity to $O((NK+L)^{2}T)$ (see Table 3).
Adaptive Interaction Mask. Due to the difference between support and query videos, we propose Adaptive Interaction Mask $\bm{J}\in{\mathbb{R}}^{(NK+L)\times(NK+L)}$ for support-support and query-support interactions, as shown in Fig. 3(a). Specifically in each task, support and query videos only aggregate semantic information from support videos while discarding interactions with query videos by element-wise multiplication of $\bm{A}$ and $\bm{J}$. By transferring knowledge between support-support and query-support videos, support features could learn intra- and inter-class correlation, and query features capture task-specific temporal cues. Besides, as shown in Fig. 3(b)(c)(d), we try other masked interaction strategies for inter-video interactions, including support-support interaction, query-support interaction and full interaction. Specifically, support-support interaction denotes ISC only considers the interactions among support videos, while query-support interaction denotes query videos only aggregate semantic information from support videos in ISC. full interaction means ISC simultaneously considers the interactions among query and support videos, but ignores the distinct characteristics of the query and support videos. The experiments about different strategies are provided in Table 4(a).
Figure 3. Different masked interaction strategies for Inter-video Semantic Correlation.
3.4. Inter-task Knowledge Transfer
Previous FSAR methods typically learn video representations within a single task, failing to explicitly learn transferable temporal knowledge between tasks. Thus, these methods lack generalizability and transferability on unseen action categories. To compensate for this limitation, we design Inter-task Knowledge Transfer (IKT) module to aggregate useful temporal knowledge from the bank, which stores diverse temporal cues shared by different tasks. IKT consists of prototype construction, knowledge retrieval, knowledge aggregation and knowledge bank updating, as shown in Fig. 2(c).
Prototype Construction. The core idea is to incorporate frame-level features into compact temporal prototypes, so as to filter out redundant information and obtain key temporal cues. We first construct $M$ learnable prototypes $\bm{P}\in{\mathbb{R}}^{M\times C}$ to summarize video representations, where we empirically set $M$ to 3 (ablation study in Table 4(c)). Given a task consisting of support and query videos, these prototypes adaptively aggregate action dynamic information from support features $\tilde{\bm{F}}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$:
(2) $\hat{\bm{P}}=\mathrm{Softmax}(\bm{P}\bm{K}_{\mathrm{t}}^{\top})\bm{V}_{\mathrm{t}}+\bm{P},$
where $\bm{K}_{\mathrm{t}}$ and $\bm{V}_{\mathrm{t}}$ are the linear transformation features of frame-level features $\tilde{\bm{F}}_{\mathrm{s}}$. As such, such prototypes $\hat{\bm{P}}$ can adaptively learn useful temporal patterns.
Knowledge Retrieval. Inspired by (Wu et al., 2018; Zhou et al., 2022), we setup a temporal knowledge bank $\bm{M}=[\bm{m}_{1},\bm{m}_{2},...,\bm{m}_{G}]\in{\mathbb{R}}^{G\times C}$ to store task-shared temporal patterns, where $G$ is the number of memory representations and $G$ is empirically set to $50$ (ablation study in Fig. 4). After obtaining original temporal prototypes $\hat{\bm{P}}$, we perform knowledge retrieval via only aggregating semantic-related memory from the bank:
(3) $\bm{P}^{\prime}=\sum_{g=1}^{G}\mathrm{topk}(\hat{\bm{P}},\bm{m}_{g})\bm{m}_{g},$
where $\mathrm{topk}$ is the function that retrieves $O$ most similar prototypes from the bank according to the semantic correspondences. We take a hyper-parameter $\kappa\!=\!O/G$ to determine the number of retrieved memory representations. $\bm{P}^{\prime}$ is retrieved historical knowledge related to current task, which could be used to further enhance task-specific video representations.
Knowledge Aggregation. Knowledge Aggregation aims to make use of learned historical knowledge to enhance temporal features. Specifically, we first deliver the union of original temporal prototypes $\hat{\bm{P}}$ and retrieved memory $\bm{P}^{\prime}$ by element-wise addition to obtain task-specific prototypes $\bar{\bm{P}}$. Then, support features $\tilde{\bm{F}}_{\mathrm{s}}$ aggregate temporal knowledge provided by prototypes $\bar{\bm{P}}$ via cross-attention mechanism:
(4) $\small\bar{\bm{F}}_{\mathrm{s}}=\texttt{CrossAttn}(\tilde{\bm{F}}_{\mathrm{s}},\bar{\bm{P}}),$
where CrossAttn is the standard cross-attention operation, and $\bar{\bm{F}}_{\mathrm{s}}$ is enhanced temporal features of $\tilde{\bm{F}}_{\mathrm{s}}$ by exploring inter-task relationships. By aggregating temporal knowledge from previous tasks, we strengthen discriminative dynamic patterns for each new episodic task.
Knowledge Bank Updating. At each episodic task, the temporal knowledge bank is continually updated to involve new learned knowledge from the current task. Our updating strategy is as follows: given $i$-th temporal prototype $\bm{p}_{i}$ in $\hat{\bm{P}}$, if the knowledge bank is not full, the prototype is directly assigned; if the bank is full, we first compute the cosine similarity between current prototype and $G$ memory representations in $\bm{M}$, and update the most similar one as follows:
(5) $\bm{m}_{t}=\texttt{SelectMax}\left(\frac{\bm{p}_{i}\cdot\bm{m}_{j}}{||\bm{p}_{i}||||\bm{m}_{j}||}\right),\bm{m}_{j}\in\bm{M},$
(6) $\bm{m}_{t}\leftarrow\mu\bm{m}_{t}+(1-\mu)\bm{p}_{i},$
where SelectMax denote the function that selects most similar memory representation from $\bm{M}$ as $\bm{m}_{t}$, and $\mu\!\in\![0,1]$ is the momentum for memory evolution. Note that during the testing phase, the temporal knowledge bank remains unchanged and does not store any information from testing data, avoiding information leakage.
More In-Depth Discussion. We emphasize that our designed Inter-task Knowledge Transfer (IKT) module strictly adheres to the standard few-shot learning protocol, and no information leakage occurs during testing. Specifically, the temporal knowledge bank in IKT only stores learned knowledge from support samples of training tasks and remains fixed during the testing phase. When performing knowledge retrieval at test time, our approach only accesses the support and query videos of the current task, and the knowledge bank used at test time remains unchanged and does not incorporate any features from test-time videos (i.e., unseen classes).
In addition, our method learns useful temporal knowledge by capturing generalizable and task-shared semantic patterns during training, rather than introducing class-specific feature leakage from testing data. This is aligned with the goal of few-shot learning, which is to generalize to new classes by utilizing prior knowledge.
3.5. Training and Testing with HR2G-shot
Training Objective. After jointly exploring three types of relation modeling, we obtain support features $\bar{\bm{F}}_{\mathrm{s}}$ and query features $\bar{\bm{F}}_{\mathrm{q}}$. Given $i$-th support frame-level features $\bar{\bm{f}}_{\mathrm{s}}^{i}\in{\mathbb{R}}^{T\times C}$ in $\bar{\bm{F}}_{\mathrm{s}}$ and $j$-th query frame-level features $\bar{\bm{f}}_{\mathrm{q}}^{j}\in{\mathbb{R}}^{T\times C}$ in $\bar{\bm{F}}_{\mathrm{q}}$, like in previous methods (Cao et al., 2020; Wang et al., 2024b), we adopt the “temporal alignment metric to obtain query-support distances:
(7) ${\mathcal{D}}=\mathtt{Metric}(\bar{\bm{f}}_{\mathrm{q}}^{j},\bar{\bm{f}}_{\mathrm{s}}^{i}),$
where $\mathtt{Metric}$ denotes the OTAM (Cao et al., 2020) metric by default. Then we can use the output support-query distances as logits to compute cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$ over the ground-truth labels.
Testing. During testing, We freeze the entire framework, and directly extracts support and query features in a single feedforward pass for the unseen test classes. Note that the temporal knowledge bank is also frozen, not incorporating any features from test-time videos. Finally, we can utilize the obtained distance in Eq. 7 as logits to produce query predictions.
Table 1. Quantitative results on SSv2-small (Goyal et al., 2017), SSv2-full (Goyal et al., 2017), HMDB51 (Kuehne et al., 2011), UCF101 (Soomro et al., 2012), and Kinetics (Carreira and Zisserman, 2017) (see §4.3). “INet-RN50” denotes ResNet-50 pre-trained on ImageNet. The best results are highlighted.
  SSv2-small SSv2-full HMDB51 UCF101 Kinetics
Method Reference Pre-training $1$-shot $5$-shot $1$-shot $5$-shot $1$-shot $5$-shot $1$-shot $5$-shot $1$-shot $5$-shot
CMN (Zhu and Yang, 2018) ECCV’18 INet-RN50 $34.4$ $43.8$ $36.2$ $48.9$ $60.5$ $78.9$
OTAM (Cao et al., 2020) CVPR’20 INet-RN50 - - $42.8$ $52.3$ $73.0$ $85.8$
AmeFuNet (Fu et al., 2020) MM’20 INet-RN50 - - - - $60.2$ $75.5$ $85.1$ $95.5$ $74.1$ $86.8$
TRX (Perrett et al., 2021) CVPR’21 INet-RN50 - $59.1$ $64.6$ $75.6$ $96.1$ $63.6$ $85.9$
TA2(Li et al., 2022) AAAI’22 INet-RN50 - - $47.6$ $61.0$ $59.7$ $73.9$ $81.9$ $95.1$ $72.8$ $85.8$
MTFAN (Wu et al., 2022) CVPR’22 INet-RN50 - - $45.7$ $60.4$ $59.0$ $74.6$ $84.8$ $95.1$ $74.6$ $87.4$
HyRSM (Wang et al., 2022) CVPR’22 INet-RN50 $40.6$ $56.1$ $54.3$ $69.0$ $60.3$ $76.0$ $83.9$ $94.7$ $73.7$ $86.1$
STRM (Thatipelli et al., 2022) CVPR’22 INet-RN50 $37.1$ $55.3$ $43.1$ $68.1$ $52.3$ $77.3$ $80.5$ $96.9$ $62.9$ $86.7$
HCL (Zheng et al., 2022) ECCV’22 INet-RN50 $38.9$ $55.4$ $47.3$ $64.9$ $59.1$ $76.3$ $82.6$ $94.5$ $73.7$ $85.8$
MoLo (Wang et al., 2023a) CVPR’23 INet-RN50 $41.9$ $56.2$ $55.0$ $69.6$ $60.8$ $77.4$ $86.0$ $95.5$ $74.0$ $85.6$
GgHM (Xing et al., 2023c) ICCV’23 INet-RN50 - - $54.5$ $69.2$ $85.2$ $96.3$ $61.2$ $76.9$ $74.9$ $87.4$
CLIP-FSAR (Wang et al., 2024b) IJCV’24 CLIP-RN50 $52.0$ $55.8$ $58.1$ $62.8$ $69.2$ $80.3$ $91.3$ $97.0$ $87.6$ $91.9$
CLIP-Freeze (Radford et al., 2021) ICML’21 CLIP-ViT-B $29.5$ $42.5$ $30.0$ $42.4$ $58.2$ $77.0$ $89.7$ $95.7$ $78.9$ $91.9$
CapFSAR (Wang et al., 2023b) Arxiv’23 CLIP-ViT-B $45.9$ $59.9$ $51.9$ $68.2$ $65.2$ $78.6$ $93.3$ $97.8$ $84.9$ $93.1$
MVP-shot (Qu et al., 2024) Arxiv’24 CLIP-ViT-B $55.4$ $62.0$ $77.0$ $88.1$ $96.8$ $99.0$ $91.0$ $95.1$
MA-FSAR (Xing et al., 2024) Arxiv’24 CLIP-ViT-B $59.1$ $64.5$ $63.3$ $72.3$ $83.4$ $87.9$ $97.2$ $99.2$ $\mathbf{95.7}$ $96.0$
CLIP-FSAR (Wang et al., 2024b) IJCV’24 CLIP-ViT-B $54.6$ $61.8$ $62.1$ $72.1$ $77.1$ $87.7$ $97.0$ $99.1$ $94.8$ $95.4$
HR2G-shot (ours) - CLIP-ViT-B $\mathbf{60.2}$ $\mathbf{66.0}$ $\mathbf{65.4}$ $\mathbf{74.8}$ $\mathbf{85.6}$ $\mathbf{88.6}$ $\mathbf{98.0}$ $\mathbf{99.3}$ 95.2 $\mathbf{96.4}$
Table 2. The impact and parameters of core components on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) under the 5-way 1-shot setting (see §4.4).
  Method Component Model Params SSv2-small HMDB51
Baseline $71.0$M $55.0$ $81.2$
[…]
IKT only $74.5$M (+3.5) $57.8$ $83.2$
[…]
4. Experiments
4.1. Experimental Setup
Dataset. We evaluate our HR2G-shot on five commonly used datasets, i.e., Kinetics (Carreira and Zisserman, 2017), SSv2-small (Goyal et al., 2017), SSv2-full (Goyal et al., 2017), HMDB51 (Kuehne et al., 2011), and UCF101 (Soomro et al., 2012). For HMDB51 and UCF101, we adopt the few-shot split from (Zhang et al., 2020; Bishay et al., 2019), with $31$/$10$/$10$ classes and $70$/$10$/$21$ classes for train/val/test, respectively. Following previous works (Wang et al., 2023a; Xia et al., 2023; Cao et al., 2022), we divide Kinetics, SSv2-small and SSv2-full into three splits: $64$/$12$/$24$ classes used for train, val, and test, respectively. The difference between SSv2-small and SSv2-full is that SSv2-full contains more videos per class in train set.
Evaluation. Following standard evaluation protocols (Cao et al., 2020; Xing et al., 2023c; Wang et al., 2023a; Liu et al., 2023; Yu et al., 2023), $5$-way $1$-shot and $5$-shot accuracy over $10,000$ tasks are used for evaluation.
4.2. Implementation Details
Network Architecture. HR2G-shot adopts pre-trained CLIP ViT-B (Radford et al., 2021) as our backbone for parameter-efficient fine-tuning, for a fair comparison with previous methods (Qu et al., 2024; Wang et al., 2024b; Xing et al., 2024). In IKT, we empirically set the size of temporal knowledge bank $G\!=\!40$ (see Fig. 4). By default, the number of temporal prototypes $M$ is set to $3$ (see Table 4(c)). For other hyper-parameters, we empirically set retrieval ratio $\kappa$ and momentum $\mu$ are to $0.8$ (see Table 4(b)) and $0.99$ (see Table 4(d)), respectively.
Network Training. In line with previous works (Cao et al., 2020; Huang et al., 2022; Wang et al., 2024b; Xing et al., 2023c), we uniformly and sparsely sample $T\!=\!8$ frames from each video to encode frame-level representations. During the training phase, random horizontal flipping and color jitter are adopted for data augmentation. We freeze both the CLIP visual encoder and text encoder, and only finetune lightweight adapters in the visual encoder as (Xing et al., 2024). We conducted training using the Adam (Kingma and Ba, 2014) optimizer with the multi-step scheduler. Due to limited space, more implementation details are left in the appendix.
Reproducibility. All experiments are conducted on three NVIDIA 3090 GPUs with 24GB memory in PyTorch. To ensure reproducibility, full code will be released.
4.3. Comparison with State-of-the-Arts
Table 1 illustrates our compelling results over the top-leading FSAR solutions on five datasets (i.e., SSv2-small (Goyal et al., 2017), SSv2-full (Goyal et al., 2017), HMDB51 (Kuehne et al., 2011), UCF101 (Soomro et al., 2012), and Kinetics (Carreira and Zisserman, 2017)). Our method reports the results on CLIP-ViT-B visual encoder. For spatial-related datasets, our method yields remarkable performance on most task settings across HMDB51, UCF101, and Kinetics. Especially for 1-shot tasks, it surpasses the previous SOTA (i.e., MA-FSAR (Xing et al., 2024)) by 2.2% on HMDB51 and 0.8% on UCF101, respectively. Temporal-related datasets (SSv2-small and SSv2-full) require models to comprehend complex temporal information, making it much more challenging. Our approach still achieves dominant results on SSv2-small and SSv2-full, surpassing other competitors across all metric task settings. This reinforces our belief that our method makes use of the temporal knowledge bank, which collects temporal cues from historical tasks, to yield discriminative temporal features. All of the above improvements across all datasets and task settings show our HR2G-shot has strong generalization for different scenes. We attribute this to the fact that we learn task-specific temporal features by transferring knowledge at both video-level and task-level.
Table 3. Ablations of different inter-video interaction manner in ISC on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) under the 5-way 1-shot setting (see §4.4).
  Method FLOPs (M) SSv2-small HMDB51
Video-level global attention $53$ $56.5$ $83.6$
Frame-level dense attention $252$ $58.9$ $84.2$
Ours $58$ $\mathbf{60.2}$ $\mathbf{85.6}$
Table 4. A set of ablation studies on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) under the $5$-way $1$-shot setting (§4.4). The adopted network designs are marked in red.
  Method SSv2-small HMDB51
support-support $58.7$ $84.7$
query-support $59.0$ $84.5$
full $57.8$ $84.0$
Ours $\mathbf{60.2}$ $\mathbf{85.6}$
(a) Masked Interaction Strategy
  Retrieval ratio $\kappa$ SSv2-small HMDB51
$\kappa=0.3$ $58.9$ $84.3$
$\kappa=0.5$ $59.5$ $84.9$
$\kappa=0.7$ $\mathbf{60.2}$ $\mathbf{85.6}$
$\kappa=0.9$ $59.9$ $85.3$
$\kappa=1.0$ $59.7$ $85.2$
(b) Retrieval Ratio
  Prototype Number $M$ SSv2-small HMDB51
$M=1$ $59.0$ $84.7$
$M=2$ $59.5$ $85.0$
$M=3$ $\mathbf{60.2}$ $\mathbf{85.6}$
$M=4$ $59.6$ $85.3$
$M=8$ $58.8$ $85.2$
(c) Prototype Number
  Momentum Coefficient $\mu$ SSv2-small HMDB51
$\mu=0$ $57.2$ $83.1$
$\mu=0.5$ $59.0$ $84.4$
$\mu=0.9$ $59.4$ $85.0$
$\mu=0.99$ $\mathbf{60.2}$ $\mathbf{85.6}$
$\mu=0.999$ $59.9$ $85.4$
(d) Momentum Coefficient
4.4. Diagnostic Experiment
Key Component Analysis. We first investigate the effectiveness of each component in HR2G-shot, i.e., Inter-video Semantic Correlation (ISC) and Inter-task Knowledge Transfer (IKT), which is summarized in Table 2. First, our proposed ISC leads to $3.7\%$ and $2.8\%$ performance gains against the baseline on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011), respectively, demonstrating the value of inter-video relation modeling. Second, after incorporating our proposed IKT into the baseline, our method improves on the two datasets by $2.8\%$ and $2.0\%$, verifying that capturing task-shared temporal knowledge by exploring inter-task relationships can yield task-specific video representations. Third, our full model HR2G-shot achieves the best performance, confirming the joint effectiveness of our overall algorithm design. Different from reported learnable model parameters in Table 5, we provide full model parameters in Table 2. As seen, our algorithm brings a modest amount of extra parameters, while leveraging such a performance leap.
Imapct of Inter-video Interaction Manner in ISC. We study the impact of our fine-grained inter-video interaction (Eq. 1) by contrasting it with video-level global attention (Wang et al., 2022) and frame-level dense attention. “video-level global attention” refers to conducting inter-video interaction via global video features, and “frame-level dense attention” means one frame token interacts with all other frame tokens in current task. As outlined in Table 3, our fine-grained inter-video interaction proves to be effective–it outperforms video-level global attention and frame-level dense attention by $3.7\%$ and $1.3\%$ on SSv2-small respectively, and efficient–its FLOPs are significantly less than frame-level dense attention. The reason why our fine-grained interaction outperforms frame-level dense attention may be that frame-level dense attention often introduces redundant or noisy interactions. Thus, we conclude that our fine-grained inter-video interaction achieves a trade-off between effectiveness and efficiency for capturing task-specific temporal cues.
Figure 4. The impact of temporal bank size $G$ in IKT on SSv2-small (Goyal et al., 2017) under the $5$-way $1$-shot setting (§4.4).
Impact of Different Masked Interaction Strategies in ISC. We further analyze the influence of our masked interaction strategy on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011). Fig. 3(b)(c)(d) shows three other alternative masked strategies, i.e., support-support interaction, query-support interaction and full interaction. As reported in Table 4(a), our masked interaction strategy (i.e., adaptive Intraction Mask) yields significant performance advancements compared with other alternatives. This suggests that designing proper masked interactions by respecting the difference between support and query videos contributes to better fine-grained cross-video interactions.
Impact of Different Retrieval Ratio $\kappa$ in IKT. We next investigate the impact of retrieval ratio $\kappa$ (Eq. 3), which controls the number of retrieved memory representations from the knowledge bank. Table 4(b) provides comparison results with regrad to different retrieval ratios under the 1-shot setting on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011). We can clearly observe that, our algorithm performs best with a relatively large retrieval ratio (i.e., $\kappa=0.7$) on SSv2-small. This verifies that we discard the semantic-unrelated temporal knowledge in the bank to alleviate the negative effects of meaningless semantic information. When $\kappa$ is too small, the performance degrades. We speculate this is because our algorithm discards too many useful memory representations related to the current task.
Figure 5. Similarity visualization between query samples ($q_{n}$) and support prototypes ($s_{n}$) with different methods in a meta-test episode from HMDB51 (Kuehne et al., 2011) (see §4.5). A higher score indicates a greater degree of similarity. The green box indicates correct prediction and the red box indicates incorrect prediction.
Impact of Temporal Prototype Number $M$ in IKT. Table 4(c) reports the performance of our HR2G-shot with regard to the number of temporal prototypes. $M\!=\!1$ means that directly treating one video as a global representation. For $M=8$, we maintain original frame-level features as temporal prototypes. After constructing learnable prototypes to summarize discriminative temporal representations, we observe consistent improvements. For example, our HR2G-shot gains the best performance (i.e., 60.2% accuracy on SSv2-small dataset) when $M\!=\!3$.
Impact of Momentum Coefficient $\mu$ in IKT. We probes the impact of momentum coefficient $\mu$ (Eq.6) on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) in Table 4(d). We can clearly observe that, our algorithm performs better with a relatively large coefficient (i.e., $\mu=0.99$), but not too large (i.e., $\mu=0.999$). When $\mu$ is too small, the performance decreases. In particular, at the extreme of no momentum (i.e., $\mu=0$), the performance significantly degrades.
Figure 6. T-SNE feature visualization of six classes learned by baseline and HR2G-shot on Kinetics (Carreira and Zisserman, 2017) and HMDB51 (Kuehne et al., 2011) ( §4.5).
[…]
Efficiency Analysis. To analyze the effectiveness of our HR2G-shot, we list comparison results with CLIP-FSAR (Wang et al., 2024b) in terms of trainable parameters, GPU memory, and inference speed in Table 5. We choose ViT-B as our visual encoder. Note that different from full model parameters in Table 2, Table 5 reports trainable parameter comparison. As shown in Table 5, compared to CLIP-FSAR, our method introduces fewer trainable parameters. Though efficient in terms of parameters, our algorithm brings slight inference delay and memory costs. However, the hierarchical relation modeling allows our method to outperform CLIP-FSAR in terms of classification accuracy, i.e., 8.5% and 5.6% accuracy improvements on HMDB51 (Kuehne et al., 2011) and SSv2-small (Goyal et al., 2017) over CLIP-FSAR, respectively.
Table 5. Complexity analysis for $5$-way $1$-shot HMDB51 (Kuehne et al., 2011) and SSv2-small (Goyal et al., 2017) evaluation. Note that there are $1$ query samples per class. Here, we report trainable parameters, GPU memory, and Speed for each model. “Acc1” and “Acc2” are the accuarcy on HMDB51 and SSv2-small, respectively (§4.4).
  Method Params Memory Speed Acc1 Acc2
CLIP-FSAR $89.5$M $14.2$G $36.7$ms $77.1$ $54.6$
HR2G-shot(ours) $19.9$M $14.3$G $39.9$ms $85.6$ $60.2$
Table 6. Generalization performance of our method with different temporal alignment metrics on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) under the $5$-way $1$-shot setting (§4.6).
  Temporal Metric SSv2-small HMDB51
CLIP-FSAR (Bi-MHM) $54.1$ $77.0$
Ours (Bi-MHM) $\mathbf{60.1}$ $\mathbf{85.0}$
CLIP-FSAR (TRX) $53.8$ $77.4$
Ours (TRX) $\mathbf{60.4}$ $\mathbf{85.7}$
CLIP-FSAR (OTAM) $54.6$ $77.1$
Ours (OTAM) $\mathbf{60.2}$ $\mathbf{85.6}$
4.5. Quality Analysis
Similarity visualization. To qualitatively demonstrate the effectiveness of hierarchical relation modeling in our HR2G-shot, we visualize the predicted similarities between query and support prototypes with different approaches (i.e., CLIP-FSAR (Wang et al., 2024b) and ours) for one task of HMDB51 in Fig. 5. As seen, our HR2G-shot can make more accurate decisions for similar classes in each task compared with CLIP-FSAR. Specifically, for the first query sample in Fig. 5, the incorrect decision obtained by CLIP-FSAR can be rectified by jointly modeling three types of relations. These results further demonstrate that our method can capture task-specific temporal cues via transferring knowledge at both video-level and task-level.
The visualization of Feature Distribution. Fig. 6 visualizes learned features of baseline and our algorithm HR2G-shot with ISC and IKT via the t-SNE tool. We clearly observe that after conducting inter-video and inter-task interactions, learned video features become more compact (intra-class) and better separated (inter-class). As shown in Fig. 6(b), the three classes “Smoke”, “Sit”, and “Pick” become clearly distinguishable from each other after conducting hierarchical relation modeling. The above phenomena suggest that we can gain task-specific temporal features by transferring knowledge at both video-level and task-level, ensuring better feature discrimination. See more examples in Appendix.
4.6. Generalization Study
To demonstrate that HR2G-shot generalizes well to different alignment metrics, Table 6 report comparison results on HMDB51 (Kuehne et al., 2011) and SSv2-small (Goyal et al., 2017) with regard to different temporal metrics (cf. Eq. 7). Different temporal metrics are adopted, including OTAM (Cao et al., 2020), TRX (Perrett et al., 2021) and Bi-MHM (Wang et al., 2022). As seen, our HR2G-shot consistently achieves better performance compared with CLIP-FSAR (Wang et al., 2024b), regardless of the temporal alignment metric employed. This shows that our method can adapt to any temporal alignment metric.
5. Conclusion
In this work, we present HR2G-shot, a novel Hierarchical Relation-augmented Representation Generalization framework for FSAR, which captures task-specific temporal cues from multiple perspectives (i.e., inter-frame, inter-video, and inter-task). Rather than only conducting inter-frame temporal interactions, we further i) perform fine-grained cross-video temporal interaction for exploring inter-video relationships, and ii) aggregate useful temporal knowledge from previous tasks for learning inter-task relationships. Our framework is hierarchical, elegant, and gains outstanding performance on five standard FSAR datasets, i.e., SSv2-full, SSv2-small, HMDB51, UCF101, and Kinetics. Our work offers a fresh perspective to current FSAR solutions by hierarchically exploring multi-level relations, and we wish it to inspire future research in this field.